Help please?!
SOLVE THE SYSTEM:
2/x+3/y-2/z=-1
3/x-18/y+5/z=4
5/x+3/y-1/z=1

Answer by DrBeeee(684)

The inverse of variables is turning you up side down! Fix it!
Merely let a = 1/x, b = 1/y and c = 1/z and yield the friendly 3x3 system of equations
2a + 3b - 2c = -1
3a - 18b + 5c = 4
5a + 3b - c = 1
Use Mathcad or your handy dandy TI-84 to get a, b and c. Then just get their inverse to get x, y, and z. You can also solve "by hand" using the inverse of a 3x3 matrix.